# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v2

@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are penalty weights.
Lemma 1.
: For a sufficiently large penalty parameter $\chi^{*}$, the solution to Problem (22) is equivalent to the solution of Problem (21).
Proof: As $\chi\rightarrow+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite. Given that the original problem’s feasible set is bounded (as argued in [24, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
To solve Problem (22) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1. Since the objective function (22a) is smooth and differentiable, the proximal step reduces to a projection. We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (19):
$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).\end{split}$ (23)
The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (24)
where
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$ (25)
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$ (26)
@@
8:   Update $\chi=\chi\times\Delta$
9:   Update $k=k+1$
10:   Update $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
11:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
12:end for
@@
Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure the convergence of the PGA method [22, 23]. In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step. The computation cost of calculating $f(\bm{\delta}_{m})$ is $\mathcal{O}(T^{2})$. Thus, the cost of calculating $\nabla f(\bm{{\delta}}_{m})$ is also $\mathcal{O}(T^{2})$. Since we have only $L_{p}$ variables, it converges in a few iterations only.
Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping as in [19, 13]. In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network. In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables. Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference from other UEs that are also on strong pilots. This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy. In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, improving performance for users on weak pilots, at the cost of higher computation in comparison to G-PFZF scheme due to projection matrix computation at the AP.
@@ III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
This projection ensures that the combiner for a weak pilot is orthogonal to the channel estimates of all strong-pilot UEs, thereby nullifying the dominant interference from that group.
@@
Furthermore
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$ (29)
These equations confirm that the projected MR (PMR) combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$. This design effectively mitigates the most significant intra-network interference, enhancing the performance of the UEs with weaker channel conditions.
By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by
@@
The derivation of the SINR expression is given in Appendix -C.
Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully decentralized manner. Here too the local weights are the same as the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -C. Accordingly, the closed-form of the local SINR or local weight is given by:
$a_{mt}=\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (31)
where
@@
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
By employing these locally computed weights, each AP independently scales its data estimates to prioritize signals with high desired gain and low interference. This process effectively embeds interference suppression functionality at the source, thereby completely eliminating the need for centralized LSFD at the CPU and the associated fronthaul overhead for channel statistics.
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is given in (31). After substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ in (31):
@@ IV Performance Evaluation
TABLE I: Baseline Simulation Parameters
Parameter Value
Number of APs ($M$) 100
Antennas per AP ($A$) 8
Number of UEs ($T$) 100
Channel bandwidth ($B$) 20 MHz
Uplink pilot symbols ($L_{p}$) 7
Coherence block length ($L_{c}$) 200 symbols
Maximum UE transmit power ($p_{\text{max}}$) 100 mW
Shadow fading standard deviation 8 dB
This section demonstrates the gains of the our proposed decentralized decoding and combining schemes in terms of the sum SE, per user SE and computation cost. We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The relevant system parameters are mentioned in Table I. The LSFCs are generated according to the model in [3], incorporating shadow fading. Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources. All APs are assumed to operate at full transmit power during both pilot training and data transmission. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, are identified using numerical simulation, to provide the best performance. To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
@@ IV-A Computation cost and Fronthaul cost
In this subsection, we analyze the computational and fronthaul signaling costs of our proposed decentralized decoding and combining schemes. The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
TABLE II: Decoding Computation and Fronthaul Overhead per UE per Coherence Block
Decoding Type Fronthaul Cost Computation Cost
LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
Proposed Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
Table II compares the fronthaul and computational costs for optimal LSFD and the proposed decentralized decoding architecture. In conventional LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block. Furthermore, LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale network deployments.
The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely. Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars. The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions required in the LSFD architecture due to centralized AP coordination. This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
Figure 2: Comparison of computational and fronthaul costs for different decoding and combining schemes.
As demonstrated in Fig. 2a, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to LSFD, particularly as the number of UEs increases. Similarly, Fig. 2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
TABLE III: Computational cost per AP per Coherence Block
@@
PFZF[19] $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AT$
G-PFZF $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AL_{p}$
PWPFZF[19] $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AT$
G-PWPFZF $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AL_{p}$
The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III. The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users. The traditional PFZF and PWPFZF schemes exhibit computational costs that scale linearly with the total number of user equipments, $T$, as in the $AT$ term for PFZF and PWPFZF. This dependence creates a significant computational bottleneck in densed networks with massive connectivity. In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs. This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population. Fig. 2c further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
IV-B Performance Evaluation

### v3

@@ III-A Generalized Partial Full-Pilot Zero-Forcing Combining Scheme
$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}\\ &+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are weights.
Lemma 2.
: For a sufficiently large penalty $\chi^{*}$, the solution to Problem (24) is equivalent to the solution of Problem (23).
Proof: As $\chi{\rightarrow}+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite. Given that the original problem’s feasible set is bounded (as argued in [23, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
For practical consideration, the penalty function should be sufficiently small (${\leq}10^{-3}$), for some value of $\chi$. In our numerical simulations, we initialize $\chi{=}1$ with $\lambda_{1}{=}5$ and $\lambda_{2}{=}10$ to ensure sufficiently small penalty function [22].
To solve Problem (24) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1. We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (21):
$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right).\end{split}$ (25)
The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\\ &-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (26)
@@
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$
@@
8:   Update $\chi=\chi\times\Delta$, $k=k+1$ and $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
9:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
10:end for
@@
III-B Generalized Protected Weak Partial Full-Pilot Zero-Forcing Combining Scheme
In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, improving performance for users on weak pilots, at the cost of higher computation in comparison to G-PFZF scheme due to projection matrix computation at the AP.
@@ III-B Generalized Protected Weak Partial Full-Pilot Zero-Forcing Combining Scheme
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$
These equations confirm that the PMR combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$. By using the combining vector (9) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by (29).
@@
The corresponding closed-form expression of d-LSFD weights for UE $t$ at $m$ for G-PWPFZF combining is:
$a_{mt}=\frac{\sqrt{p^{{u}}_{t}}b_{tt}^{m}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}},\ \ \ \text{where},$ (30)
@@
${W}_{{m}t}=\sum\nolimits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
The derivation of the closed-form SINR and d-LSFD expressions are given in Appendix -E.
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is $\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{S_{mt}}{I_{mt}}$, where
@@
Convergence and Computation Cost Analysis of PGA: Since the feasible set defined by the box constraint (21) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -F) with constant L>0L{>}0, one should choose $\alpha{\in}(0,1/L]$ to ensure the convergence of the PGA method [21, 22]. Step size $\alpha$ is selected using backtracking line-search to achieve convergence for Algorithm 1.
At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step and is given by $\mathcal{O}(TL_{p})$. For both proposed schemes, the average number of PGA iterations per AP is approximately 7, regardless of whether the number of UEs is 50 or 100. This indicates that the convergence speed of PGA is largely insensitive to UE density. The average runtime per AP on a standard laptop equipped with an Intel Core Ultra 7 processor and 16 GB RAM is approximately 0.01 s for 50 UEs and increases to about 0.02–0.03 s for 100 UEs. This increase is mainly due to the higher cost of objective evaluation rather than slower convergence. These results confirm that the proposed PGA-based pilot grouping is computationally efficient and scalable to dense UE deployments.
@@ IV Performance Evaluation
We consider a network with $M=100$ APs with $A=8$ antennas each and $T=100$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The LSFCs are generated according to the model in [3], incorporating shadow fading with standard deviation of 8 dB. Pilots are assigned using the scalable pilot assignment strategy from [3]. All UEs are assumed to operate at full transmit power of 100 mW. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using $90\%$ of total LSFCs received at the AP, as identified via numerical simulation, to provide the best performance. Also $\mathcal{K}_{t}$ denotes the set of UEs that have atleast one serving AP in common with UE $t$. To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
@@ IV-A Computation cost and Fronthaul cost
We analyze the computational and fronthaul signaling costs of our proposed d-LSFD and combining schemes. The costs are computed following the approach outlined in [17, 18, 12, 8], where only multiplication and division operations are counted, as these dominate the computational load.
TABLE I: Decoding Computation and Fronthaul Overhead per Coherence Block
Decoding Fronthaul Computation Cost
c-LSFD [17] $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|+2TM$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}\frac{|\mathcal{P}_{i_{t}}|+2}{2}+\sum_{t=1}^{T}|\mathcal{M}_{t}|T+\sum_{t=1}^{T}\frac{|\mathcal{M}_{t}|^{3}-|\mathcal{M}_{t}|}{3}$
p-LSFD [17] $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|+2TM$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}\frac{|\mathcal{P}_{i_{t}}|+2}{2}+\sum_{t=1}^{T}|\mathcal{M}_{t}||\mathcal{K}_{t}|+\sum_{t=1}^{T}\frac{|\mathcal{M}_{t}|^{3}-|\mathcal{M}_{t}|}{3}$
d-LSFD $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|T+\sum_{t=1}^{T}|\mathcal{M}_{t}|+\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}$
Table I compares the fronthaul and computational costs for c-LSFD, p-LSFD and proposed d-LSFD. In c-LSFD and p-LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead. Furthermore, c-LSFD and p-LSFD require computationally expensive matrix operations such as matrix inversion at the CPU. The d-LSFD eliminates these bottlenecks by avoiding centralized decoding entirely, reducing the fronthaul load for sending data estimates only. The computational cost is similarly streamlined by eliminating the matrix inversions required in c-LSFD.
Figure 2: Comparison of computational and fronthaul costs for different decoding schemes.
As demonstrated in Fig. 2a, the proposed d-LSFD weights computation provides substantial computation cost reduction compared to c-LSFD, particularly as the number of UEs increases. Similarly, Fig. 2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
TABLE II: Computational cost per AP per Coherence Block
@@
PFZF[18] $\underbrace{\frac{3L_{\mathcal{S}_{m}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{m}}A}{2}{+}\frac{L_{\mathcal{S}_{m}}^{3}{-}L_{\mathcal{S}_{m}}}{3}}_{C_{m}}{+}A|\mathcal{T}_{m}|$
G-PFZF $C_{m}{+}AL_{p}$
PWPFZF[18] $C_{m}{+}2(L_{p}{-}L_{\mathcal{S}_{m}})L_{\mathcal{S}_{m}}A{+}A|\mathcal{T}_{m}|$
G-PWPFZF $C_{m}{+}2(L_{p}{-}L_{\mathcal{S}_{m}})L_{\mathcal{S}_{m}}A{+}AL_{p}$
The PFZF and PWPFZF schemes exhibit costs that scale linearly with the total number of serving UEs, $|\mathcal{T}_{m}|$, as shown in Table II. In contrast, the proposed G-PFZF and G-PWPFZF schemes cost scales only with the number of pilots, $L_{p}$.
IV-B Numerical Simulations
Figure 3: Empirical distribution versus SINR lower bound in Theorem 2 across different network realizations.
First, we empirically evaluate the tightness of the bound presented in Theorem 2. We also numerically verify Proposition 1 and demonstrate that, for ZF/MR-type combining, the bound is tight and depends on the level of pilot contamination.
Fig. 3 shows the empirical distribution of the bound $\max(0,1-\kappa_{t}\rho_{t}^{2})$ for pilot lengths $L_{p}\in{7,15,25,T}$ and for different local combining schemes. For ZF and MR combining, increasing the pilot length shifts the distribution toward the interval 0.80.811, indicating that the d-LSFD solution closely approximates the c-LSFD in most realizations. When $L_{p}=T$, the bound becomes exactly one for both ZF and MR combining, consistent with Proposition 1. For these schemes, the off-diagonal entries of $\mathbf{J}_{t}$ are dependent on pilot contamination, which decreases with increase in pilot length, allowing the d-LSFD weights to remain well aligned with the centralized solution.
In contrast, for the MMSE scheme, the bound is concentrated near the interval 00.20.2 even when pilot contamination is removed. This behavior occurs because the MMSE combining vector depends on the channel estimates of all UEs, regardless of whether they share the same pilot sequence. As a result, the off-diagonal entries of the interference matrix $\mathbf{J}_{t}$ remain significant, and discarding them in the d-LSFD design leads to large deviations from the centralized solution. These results confirm that the proposed d-LSFD scheme is highly effective for ZF/MR-type combining schemes but unsuitable for combiners with strong global coupling, such as MMSE.